The seating capacity of a hockey arena is 18,000. With the ticket price set at $100, attendance at recent games has been 12,000. A market survey indicates that for every dollar the ticket price is lowered, the average attendance will increase by 130.
Write a quadratic formula to indicate this change.

Answer by lwsshak3(11628)

The seating capacity of a hockey arena is 18,000. With the ticket price set at $100, attendance at recent games has been 12,000. A market survey indicates that for every dollar the ticket price is lowered, the average attendance will increase by 130.
Write a quadratic formula to indicate this change.
**
let x=number of $1 reductions in ticket price
100-x=reduced price
130x=increase in attendance
12000+130x=new attendance
revenue=price*attendance
revenue=(100-x)(12000+130x)